# RED-DiffEq: Regularization by denoising diffusion models for solving inverse PDE problems with application to full waveform inversion

### v1

@@ 2.3 Validation on the OpenFWI benchmark
We conducted a comprehensive evaluation of RED-DiffEq on the OpenFWI benchmark [62] under three distinct scenarios: clean seismic data, seismic data with Gaussian noise contamination, and seismic data with missing traces. The OpenFWI dataset has a collection of numerical solutions to the acoustic wave equation representing diverse geological structures, providing a comprehensive benchmark for evaluating seismic inversion methods. For our experiments, we selected four particularly challenging families from this dataset: Curve Fault (CF-B), Flat Velocity (FV-B), Flat Fault (FF-B), and Curve Velocity (CV-B) (see some examples in Fig. 2c). Each family represents distinct geological structures that are representative of complex subsurface formations encountered in real-world scenarios. Specifically, CF-B features intricate curved fault systems with varying angles and intersections. FV-B consists of horizontally layered velocity structures. FF-B contains straight fault lines cutting through the model. Lastly, CV-B exhibits curved velocity variations with smooth transitions between layers. For benchmarking, we compared our method against established physics-driven approaches, including standard FWI (no regularization), FWI with Tikhonov regularization, FWI with total variation (TV) regularization, and a recent diffusion model based method [58].
@@ 2.3.1 Clean seismic data
We first evaluated various methods on clean seismic data using 100 previously unseen samples from each of the four chosen geological families in the OpenFWI dataset, and all the quantitative metrics are the average of these 400 test cases. Each sample was optimized for 300 iterations, starting from an initial model obtained by applying a Gaussian filter (with a standard deviation $\sigma=10$) to the ground-truth velocity model. Fig. 4a shows representative velocity inversion examples across the four different geological families. The unregularized standard FWI method produces a large error and fails to capture essential geological features. While Tikhonov regularization improves upon the standard method, it produces an overly smoothed model that loses critical details at layer boundaries. TV preserves sharp edges but introduces “staircase” artifacts that distort the geology. The diffusion-based approach by Wang et al. shows promise with simple structures but struggles with fine-scale fault structures. Our RED‑DiffEq framework outperforms all baselines across the four geological families, simultaneously preserving high‑contrast velocity discontinuities, such as fault planes, while maintaining the smooth lateral and vertical velocity gradients that characterize individual strata. RED‑DiffEq produces the most faithful reconstructions, even for the complex CV‑B models (Fig. 4a, last row).
Figure 4: Results on the OpenFWI dataset. (a)Qualitative comparison among different regularization methods. (b) Quantitative comparison among different regularization methods through 300 iterations. (c) Vertical profile results of different regularization methods. (d) Performance of different regularization methods on different noise levels. (e) Performance of different regularization methods on different number of missing traces.
The convergence analysis (Fig. 4b, top row) reveals that RED-DiffEq consistently achieves the lowest RMSE and MAE, and highest SSIM throughout the optimization process. Statistical analysis using raincloud plots (Fig. 4b, bottom row) further confirms RED-DiffEq’s superior performance, revealed by narrower error distributions centered on better median values than other methods. Moreover, the more compact interquartile range indicates greater consistency and stability across diverse geological scenarios.
@@ 2.3.2 Noisy seismic data
We evaluated the robustness of various inversion methods by adding Gaussian noise with standard deviations from 0.1 to 0.5 to the seismic data. The comparative results reveal distinct behaviors across different approaches under noisy conditions (Fig. 4d).
RED-DiffEq exhibits superior noise resilience, maintaining sharp geological boundaries and coherent structural features at all noise levels. Most notably, it accurately reconstructs the distinct layered structures and preserves the continuous interfaces that are evident in the ground truth (Fig. 4d, right column). In contrast, the unregularized FWI approach shows significant noise-induced artifacts, particularly in the deeper regions of the velocity model, with increasing deterioration at higher noise levels. Tikhonov regularization reduces noise but introduces excessive smoothing, blurring the sharp boundaries between layers present in the ground truth. While total variation regularization preserves some sharp boundaries, it produces notable staircase artifacts that erase finer details, especially when the standard deviation of noise is 0.5, which further hinders the geological interpretability of the result. The method by Wang et al. shows moderate improvement over traditional approaches but struggles to maintain structural coherence under high noise conditions, exhibiting patchy artifacts in the reconstructed velocity model. By comparison, RED-DiffEq achieves the best balance between noise suppression and feature preservation, maintaining geological plausibility and interpretability even under severe noise contamination.
@@ 2.3.3 Seismic data with missing traces
We evaluated the performance of all methods in missing trace scenarios by randomly removing 5 to 20 traces from the seismic data. All methods exhibit solid robustness (Fig. 4e), with only gradual performance degradation as the number of missing traces increases. This behavior highlights the physics-driven resilience of FWI, as further supported by the two-way ANOVA test in Section S5, where the main effect of missing traces is not statistically significant ($p>0.05$).
However, visual examination of the reconstructed velocity models reveals notable differences in reconstruction quality (Fig. 4e, right column). RED-DiffEq maintains the highest fidelity of geological structures, successfully preserving all the essential features in the upper and bottom regions of the velocity model. In contrast, all other methods, while stable in metrics, shows visible deterioration in resolving deeper structures as missing traces increase.
These results highlight an important distinction between quantitative metrics and qualitative geological interpretability in FWI. While missing traces may not significantly impact numerical accuracy, the ability to maintain geologically meaningful reconstructions becomes a crucial differentiator among methods.
2.4 Validation on the Marmousi benchmark
To further evaluate our method’s performance, we conducted experiments on the widely used Marmousi model [63], which features complex geological structures, sharp velocity contrasts, and intricate subsurface heterogeneity. This challenging benchmark differs substantially from the OpenFWI dataset used to train our diffusion model, thus providing a rigorous test of out-of-distribution performance.
@@ 2.4 Validation on the Marmousi benchmark
We initialized the velocity model by applying a Gaussian filter (with $\sigma$ ranging from 20 to 30) to the ground-truth model, effectively removing fine-scale details to mimic practical scenarios characterized by limited prior information. RED-DiffEq accurately reconstructs critical geological features within 300 iterations, recovering complex fault systems, sharp velocity discontinuities, and layered strata (Fig. 5a). Crucially, it maintains this high-fidelity performance across all considered initial conditions, as evidenced by the consistently superior results in RMSE, MAE, and SSIM (Fig. 5b). While these metrics do show a mild sensitivity to initialization at higher $\sigma$, the degradation remains minimal compared to that of benchmark methods. In contrast, unregularized physics-driven methods exhibit high-frequency artifacts, Tikhonov regularization yields only marginal improvements, TV regularization suffers from staircase effects, and Wang et al.’s diffusion-based technique shows limited generalization capacity, performing even worse than the physics-driven baseline in these more complex settings.
@@
Figure 5: Results on the Marmousi model. (a) Qualitative comparisons among different regularization methods. (b) Quantitative performance of each method under various initial parameters. The variable $\sigma$ indicates the standard deviation of the Gaussian filter applied to the ground-truth data to generate the initial model. Larger $\sigma$ implies less preserved information. (c) Vertical velocity profiles at 4.9 km position comparing the inversion results to the ground truth.
@@
To examine velocity reconstructions in greater detail, we compare vertical profiles at a position of 4.9 km for Marmousi (Fig. 5c). RED-DiffEq closely follows the ground truth across diverse depth intervals, capturing both sharp velocity transitions and more gradual gradients. Other methods deviate substantially at mid-range to deeper sections, underscoring their difficulties in preserving critical geological features under limited initial information.
2.5 Validation on the Overthrust benchmark
@@
We extended our evaluation to the Overthrust model [64], which also presents complex geological structures, sharp velocity contrasts, and intricate subsurface heterogeneity. This model serves as another challenging benchmark that differs from our training data, further testing our method’s generalization capabilities.
@@
Figure 6: Results on the Overthrust model. (a) Qualitative comparisons among different regularization methods. (b) Quantitative performance (RMSE, MAE, and SSIM) of each method under various initial parameters. The variable $\sigma$ indicates the standard deviation of the Gaussian filter applied to the ground-truth data to generate the initial model. Larger $\sigma$ implies less preserved information. (c) Vertical velocity profiles at 3.5 km position comparing the inversion results to the ground truth.
The vertical velocity profile at 3.5 km for the Overthrust model (Fig. 6c) demonstrates RED-DiffEq’s ability to accurately capture velocity variations across different depths. Other methods struggle with accurately recovering the velocity structure, particularly in deeper sections. RED-DiffEq’s performance on the Overthrust model further confirms its ability to generalize to different geological settings while maintaining high reconstruction fidelity, demonstrating its practical value for real-world seismic imaging applications.
@@ 3 Conclusions
We introduce RED-DiffEq, a framework that integrates diffusion models as a regularization method for solving inverse PDE problems. Through comprehensive experiments on FWI, we have demonstrated the key advantages of RED-DiffEq. RED-DiffEq not only performs better than other methods on clean seismic data, but also shows remarkable robustness against two types of data imperfections: Gaussian noise contamination and missing traces, where conventional methods typically fail or require careful parameter tuning. RED-DiffEq also demonstrates strong generalization capabilities beyond its training data distribution. Despite being trained only on the OpenFWI dataset, it successfully reconstructs complex geological structures in the challenging Marmousi and Overthrust benchmarks.
The RED-DiffEq framework is inherently general and could be easily adapted to other PDE-governed inverse problems. RED-DiffEq represents a step forward in combining data-driven deep learning with physics-based inversion, offering a promising direction for solving challenging inverse problems while maintaining robust performance and physical consistency. Future research directions include incorporating physics-informed constraints in the diffusion model training and exploring applications to other inverse problems.
@@ 4.1 RED-DiffEq framework
We present RED-DiffEq as a two-stage framework that integrates a learned prior into a PDE-governed inverse problem via regularization by denoising (RED). The method comprises two stages: (1) a pretraining stage that learns a generative prior and provides an approximation to the marginal score of the noisy data distribution (Fig. 2), and (2) an inversion stage that uses a Tweedie-inspired denoise operator as a regularizer (Fig. 3). In the RED regularizer (Eq. (2)), the operator $\mathcal{D}_{\theta}$ encourages the current solution $\mathbf{x}_{k}$ toward regions of higher plausibility under the prior distribution with density $p_{\text{data}}$.
@@
$q\!\big(\mathbf{x}_{\text{data},t}\mid\mathbf{x}_{\text{data}}\big)\;=\;\mathcal{N}\!\Big(\mathbf{x}_{\text{data},t};\ \sqrt{\gamma(t)}\,\mathbf{x}_{\text{data}},\ (1-\gamma(t))\,\mathbf{I}\Big),$
@@
$\mathcal{L}_{\text{DDPM}}(\theta)\;=\;\mathbb{E}_{\,t\sim\mathcal{U}\{1:T\},\,\mathbf{x}_{\text{data}}\sim P_{\text{data}},\,\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\!\left[\left\|\,\boldsymbol{\epsilon}-\hat{\boldsymbol{\epsilon}}_{\theta}\!\big(\mathbf{x}_{\text{data},t},\,t\big)\,\right\|_{2}^{2}\right].$
@@
During experiments, we use a single-sample Monte Carlo estimate of the expectation. Moreover, we empirically find that dropping $w(t)$ yields better convergence and performance. Therefore, we fix the regularization weight empirically as a constant $\lambda$.
@@
We minimize Eq. (13) with respect to $\mathbf{x}_{k}$ using gradient descent optimization. To reduce the computational cost, we stop gradients through the denoiser for computing $\nabla_{\mathbf{x}}R(\mathbf{x}_{k})$, i.e., we treat $\hat{\boldsymbol{\epsilon}}_{\theta}$ as a constant with respect to $\mathbf{x}_{k}$ in the regularizer term, whose performance has been empirically validated in Refs. [57, 66]. At each iteration, we resample $(t,\boldsymbol{\epsilon})$, which introduces stochasticity to the optimization.

### v2

@@ 2.3 Validation on the OpenFWI benchmark
We conducted a comprehensive evaluation of RED-DiffEq on the OpenFWI benchmark [12] under three distinct scenarios: clean seismic data, seismic data with Gaussian noise contamination, and seismic data with missing traces. The OpenFWI dataset comprises pairs of velocity models and simulated seismic wavefields representing diverse geological structures, providing a comprehensive benchmark for evaluating seismic inversion methods. For our experiments, we selected four particularly challenging families from this dataset: Curve Fault (CF-B), Flat Velocity (FV-B), Flat Fault (FF-B), and Curve Velocity (CV-B) (see some samples in Fig. 2c). Each family exhibits distinct geological structures that are representative of complex subsurface formations encountered in real-world scenarios. Specifically, CF-B features intricate curved fault systems with varying angles and intersections. FV-B consists of horizontally layered velocity structures. FF-B contains straight fault lines cutting through the model. Lastly, CV-B exhibits curved velocity variations with smooth transitions between layers. For benchmarking, we compared our method against established physics-driven approaches, including standard FWI (no regularization), FWI with Tikhonov regularization, FWI with total variation (TV) regularization, and recent diffusion-model-based methods for FWI [58, 54].
@@ 2.3.1 Clean seismic data
We first evaluated various methods on clean seismic data using 100 previously unseen samples from each of the four chosen geological families in the OpenFWI dataset. All the quantitative metrics are the average of these 400 test cases. For optimization-based approaches (RED-DiffEq, Tikhonov (Sec. S4.6), and TV (Sec. S4.7)), each sample was optimized for 300 iterations. For the diffusion-sampling-based baselines (DiffusionFWI and its variant with the ILVR technique, namely DiffusionILVR), we adopted the protocol suggested in the original implementation [58], initializing the reverse diffusion process at $t=100$ with 10 conventional FWI gradient steps per diffusion timestep; please refer to Sec. S4.2 for more details about hyper-parameters and design choice, and Sec. S4.4 for details about the forward solver used for FWI. All the methods started from an initial model obtained by applying a Gaussian filter (with a standard deviation $\sigma=10$) to the ground-truth velocity model. Fig. 4a shows representative velocity inversion examples across the four different geological families. The unregularized standard FWI method produces a large error and fails to capture essential geological features. While Tikhonov regularization improves upon the standard method, it produces an overly smoothed model that loses critical details at layer boundaries. TV preserves sharp edges but introduces “staircase” artifacts that distort the geology. DiffusionFWI and DiffusionILVR show promise with simple structures but struggle with fine-scale structures. Our RED‑DiffEq framework outperforms all baselines across the four geological families, simultaneously preserving high‑contrast velocity discontinuities, such as fault planes, while maintaining the smooth lateral and vertical velocity gradients that characterize individual strata. RED‑DiffEq produces the most faithful reconstructions, even for the complex CV‑B models (Fig. 4a, last row).
Figure 4: Results on the OpenFWI dataset. (a) Qualitative comparison among different methods. (b) Quantitative comparison among different methods. (c) Vertical profile results of different methods. (d) Performance of different methods under different Gaussian noise levels in the seismic data. (e) Performance of different methods under different Laplacian noise levels in the seismic data. (f) Performance of different methods under different number of missing traces in the seismic data.
The convergence analysis (Fig. 4b, top row) reveals that RED-DiffEq consistently achieves the lowest root mean squared error (RMSE) and mean absolute error (MAE), and highest structural similarity index measure (SSIM) throughout the optimization process. We note that DiffusionFWI and DiffusionILVR utilize a nested generation process distinct from standard optimization iterations, rendering a direct convergence curve comparison inapplicable. We compare the final inversion performance across all methods (Fig. 4b, bottom row), where RED-DiffEq consistently outperforms all benchmarks.
@@ 2.3.2 Noisy seismic data
We evaluated the robustness of various inversion methods by adding Gaussian noise and Laplacian noise with standard deviations and scales ranging from 0.1 to 0.5 to the seismic data. Specifically, Gaussian noise levels (standard deviations ranging from 0.1 to 0.5) correspond to an average Signal-to-Noise Ratio (SNR) range of 24.21 dB to 10.23 dB. Similarly, Laplacian noise levels (scales ranging from 0.1 to 0.5) correspond to an SNR range of 21.20 dB to 7.23 dB. In scientific inversion, an SNR that is around or less than 10 dB is considered challenging in general. The comparative results reveal distinct behaviors across different approaches under different noisy conditions (Figs. 4d and e).
All diffusion-model-based methods exhibit noise resilience, and RED-DiffEq more accurately reconstructs distinct layered structures and preserves the continuous interfaces that are evident in the ground truth. In contrast, the unregularized FWI approach shows significant noise-induced artifacts, particularly in the deeper regions of the velocity model, with increasing deterioration at higher noise levels. Tikhonov regularization reduces the impact of measurement noise but introduces excessive smoothing, blurring the sharp boundaries between layers present in the ground truth. While total variation regularization preserves some sharp boundaries, it produces notable staircase artifacts that erase finer details, especially when the noise level reaches 0.5, which further hinders the geological interpretability of the result.
@@ 2.3.3 Seismic data with missing traces
We evaluated the performance of all inversion methods in missing-trace scenarios by randomly removing 15 to 60 traces from the seismic data. Since each source location (shot gather) contains 70 traces in total, this corresponds to losing between 21.4% and 85.7% of the available acquisition information. Crucially, the specific indices of the missing traces were held constant across all source gathers for a given velocity model. That is, the set of active receiver locations remained identical for every shot in the experiment.
As the proportion of missing data increases, all methods experience performance degradation (Fig. 4f). However, diffusion-model-based methods exhibit significantly stronger robustness to sparse acquisition compared to standard baselines. RED-DiffEq consistently achieves the lowest RMSE and MAE and the highest SSIM across all missing-trace ratios, demonstrating superior stability under incomplete seismic measurements.
Qualitative comparisons in Fig. 4f further confirm these observations. When a moderate number of traces are missing (15 traces), RED-DiffEq successfully reconstructs layered structures and continuous interfaces that closely resemble the ground truth. Even in the extreme case of 60 missing traces, RED-DiffEq preserves coherent subsurface structures and suppresses severe artifacts, whereas most benchmark methods fail to recover meaningful geological features. In contrast, the unregularized FWI approach suffers from strong instability, while Tikhonov and Total Variation regularization methods produce overly smooth or staircase-like reconstructions that obscure fine-scale details. These results demonstrate that RED-DiffEq effectively mitigates the ill-posedness introduced by sparse acquisition, providing reliable reconstructions even under severely incomplete seismic data.
2.4 Uncertainty quantification of RED-DiffEq
We evaluate the uncertainty quantification (UQ) capability of RED-DiffEq on the OpenFWI benchmark. For each seismic sample, we generated an ensemble of 20 reconstructions by executing the RED-DiffEq inversion with different random seeds, explicitly leveraging the stochastic sampling inherent in the diffusion-based regularizer. We adopt the ensemble mean as the final predicted velocity model and utilize the pixel-wise standard deviation as the metric for spatial uncertainty.
The results in Fig. 5 indicate that high uncertainty is concentrated in regions of structural complexity (e.g., fault lines), and this predicted uncertainty exhibits a positive correlation with the absolute reconstruction error measured between the ground truth and the ensemble mean of the reconstructed velocity models. This relationship is further corroborated by the middle panel and vertical profiles in Fig. 5. Across all 400 OpenFWI test models, the average Spearman rank ($R$) correlation between absolute error and predicted uncertainty is 0.66, and the average Pearson correlation ($\rho$) is 0.52. Furthermore, the 95% confidence intervals derived from the ensemble consistently contain the ground-truth profiles for the selected samples, suggesting that the uncertainty estimates are well-calibrated.
Figure 5: Examples of uncertainty quantification of RED-DiffEq. The ensemble mean and standard deviation are computed using an ensemble size of 20. We illustrate the correlation between pixel-wise absolute error and predicted uncertainty, where $R$ represents the Spearman rank, and $\rho$ represents the Pearson correlation. Vertical profiles display the mean prediction with 95% confidence intervals at the midpoint (350 m) of each velocity model.
2.5 Validation on the Marmousi and Overthrust benchmarks
To further evaluate our method’s performance, we conducted experiments on the widely used Marmousi model [56]. This challenging benchmark differs substantially from the OpenFWI dataset used to train our diffusion model. While OpenFWI primarily consists of canonical velocity fields with simplified fault geometries and moderate lateral variability, the Marmousi model simulates a deltaic sedimentary environment characterized by rapid sediment accumulation, salt tectonics (salt creep), and growth faults. This provides a test of out-of-distribution performance and domain decomposition capacity on a more complex synthetic dataset with a larger domain.
@@ 2.5 Validation on the Marmousi and Overthrust benchmarks
We initialized the velocity model by applying a Gaussian filter (with $\sigma$ ranging from 20 to 30) to the ground-truth model, effectively removing fine-scale details to mimic practical scenarios characterized by limited prior information. RED-DiffEq accurately reconstructs critical geological features within 300 iterations, recovering complex fault systems, sharp velocity discontinuities, and layered strata (Fig. 6a). Crucially, it maintains this high-fidelity performance across all considered initial conditions, as evidenced by the consistently superior results in RMSE, MAE, and SSIM (Fig. 6b top row). We note that the baselines DiffusionFWI and DiffusionILVR use the patch-based sliding window approach described in the original paper [58], which differs from the DDR strategy utilized in RED-DiffEq.
We further evaluated robustness against Gaussian and Laplacian noise in the seismic data where the initial velocity model is obtained by applying a Gaussian filter with $\sigma=20$. For Gaussian noise with standard deviations ranging from 0.1 to 0.5, the corresponding SNR decreases from 18.42 dB to 4.44 dB. For Laplacian noise with scales ranging from 0.1 to 0.5, the corresponding SNR decreases from 15.41 dB to 1.43 dB. While all methods exhibit reasonable stability in terms of RMSE and MAE (Fig. 6b, second and third rows), their performance degrades dramatically in terms of SSIM. Notably, while DiffusionFWI and DiffusionILVR exhibit strong robustness against noise, RED-DiffEq still outperforms all methods across noise levels from 0.1 to 0.5.
@@
Figure 6: Results on the Marmousi model. (a) Qualitative comparisons among different methods with initial model $\sigma=20$. (b) Quantitative performance of each method under various initial parameters $\sigma$ (first row), Gaussian noise levels (second row), and Laplacian noise levels (third row). The variable $\sigma$ indicates the standard deviation of the Gaussian filter applied to the ground-truth data to generate the initial model. Larger $\sigma$ implies less preserved information. (c) Vertical velocity profiles at the midpoint position comparing the inversion results to the ground truth.
@@
To examine velocity reconstructions in greater detail, we compare vertical profiles at midpoint for Marmousi, where results are derived from an initial velocity model generated by applying a Gaussian filter with $\sigma=20$ (Fig. 6c). RED-DiffEq closely follows the ground truth across diverse depth intervals, capturing both sharp velocity transitions and more gradual gradients. Other methods deviate substantially at mid-range to deeper sections, underscoring their difficulties in preserving critical geological features under limited initial information.
@@
We extended our evaluation to the Overthrust model [2], which serves as another challenging benchmark that differs from our training data, further testing our method’s generalization and domain decomposition capabilities. Specifically, the Overthrust model simulates a complex onshore geological environment featuring large-scale thrust faults, repeated stratigraphic units, and discontinuous velocity fields that are absent in more canonical and simpler velocity model distributions of OpenFWI.
Figure 7: Results on the Overthrust model. (a) Qualitative comparisons among different methods with initial model $\sigma=20$. (b) Quantitative performance (RMSE, MAE, and SSIM) of each method under various initial parameters $\sigma$ (first row), Gaussian noise levels (second row), and Laplacian noise levels (third row). The variable $\sigma$ indicates the standard deviation of the Gaussian filter applied to the ground-truth data to generate the initial model. Larger $\sigma$ implies less preserved information. (c) Vertical velocity profiles at the midpoint position comparing the inversion results to the ground truth.
@@
Similarly, we tested the robustness against Gaussian and Laplacian noise for all methods where results are obtained with an initial velocity model by applying a Gaussian filter with $\sigma=20$. For Gaussian noise with standard deviations ranging from 0.1 to 0.5, the corresponding SNR decreases from 19.27 dB to 5.28 dB. For Laplacian noise with scales ranging from 0.1 to 0.5, the corresponding SNR decreases from 16.25 dB to 2.28 dB. Unlike the Marmousi case, conventional regularization methods suffer significantly more from noise contamination in this setting, evidenced by worse RMSE, MAE, and SSIM scores. DiffusionFWI and DiffusionILVR continue to exhibit very strong robustness; however, RED-DiffEq still outperforms other methods under all considered scenarios.
The vertical velocity profile at midpoint for the Overthrust model, where results are obtained with an initial velocity model by applying a Gaussian filter with $\sigma=20$ (Fig. 7c), demonstrates RED-DiffEq’s ability to accurately capture velocity variations across different depths. Other methods struggle with accurately recovering the velocity structure, particularly in deeper sections. RED-DiffEq’s performance on the Overthrust model further confirms its ability to generalize to different geological settings while maintaining high reconstruction fidelity, demonstrating its practical value for real-world seismic imaging applications.
@@ 3 Conclusions
We introduce RED-DiffEq, a framework that integrates diffusion models as a regularization method for solving inverse PDE problems, with a particular focus on full waveform inversion (FWI). Through comprehensive experiments, we have demonstrated the key advantages of RED-DiffEq. RED-DiffEq not only performs better than other methods on clean seismic data, but also shows remarkable robustness against three types of data imperfections: Gaussian and Laplacian noise contamination and missing traces in the seismic data, where conventional methods typically fail or require careful parameter tuning. In addition, RED-DiffEq shows the capacity for producing meaningful uncertainty quantification. We note, however, that the calibration quality of ensemble-based uncertainty estimates may degrade in out-of-distribution scenarios, where the training data does not fully capture the target geological complexity. RED-DiffEq also demonstrates strong generalization and domain decomposition capabilities beyond its training data distribution and domain size. Despite being trained only on the OpenFWI dataset, it successfully reconstructs complex geological structures with a larger domain size in the challenging Marmousi and Overthrust benchmarks. We note that the experiments in this work are conducted on synthetic benchmarks. Applying RED-DiffEq to real field seismic data requires access to proprietary datasets and careful adaptation to field-specific noise characteristics, acquisition geometries, and significantly larger spatial scales. Addressing these challenges and validating our approach on real field data constitute important directions for future work.
@@ 4.1 RED-DiffEq framework
We present RED-DiffEq as a two-stage framework that integrates a learned prior into a PDE-governed inverse problem via regularization by denoising (RED). The method comprises two stages: (1) a pretraining stage that learns a generative prior and provides an approximation to the marginal score of the noisy data distribution (Fig. 2 and Sec. S4.1), and (2) an inversion stage that uses a Tweedie-inspired denoise operator as a regularizer (Fig. 3). In the RED regularizer (Eq. (3)), the operator $\mathcal{D}_{\theta}$ encourages the current solution $\mathbf{x}_{k}$ toward regions of higher plausibility under the prior distribution with density $p_{\text{data}}$.
@@
$q\!\big(\mathbf{x}_{\text{data},t}\mid\mathbf{x}_{\text{data}}\big)\;=\;\mathcal{N}\!\Big(\mathbf{x}_{\text{data},t};\ \sqrt{\gamma(t)}\,\mathbf{x}_{\text{data}},\ (1-\gamma(t))\,\mathbf{I}\Big),$ (11)
@@
$\mathcal{L}_{\text{DDPM}}(\theta)\;=\;\mathbb{E}_{\,t\sim\mathcal{U}\{1:T\},\,\mathbf{x}_{\text{data}}\sim P_{\text{data}},\,\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\!\left[\left\|\,\boldsymbol{\epsilon}-\hat{\boldsymbol{\epsilon}}_{\theta}\!\big(\mathbf{x}_{\text{data},t},\,t\big)\,\right\|_{2}^{2}\right].$ (13)
@@
During experiments, we use a single-sample Monte Carlo estimate of the expectation. Moreover, we empirically find that dropping $w(t)$ yields better convergence and performance (Sec. S4.3). Therefore, we fix the regularization weight empirically as a constant $\lambda$.
@@
We minimize Eq. (18) with respect to $\mathbf{x}_{k}$ using gradient descent optimization. To reduce the computational cost, we stop gradients through the denoiser for computing $\nabla_{\mathbf{x}}R(\mathbf{x}_{k})$, i.e., we treat $\hat{\boldsymbol{\epsilon}}_{\theta}$ as a constant with respect to $\mathbf{x}_{k}$ in the regularizer term, whose performance has been empirically validated in Refs. [39, 43]. At each iteration, we resample $(t,\boldsymbol{\epsilon})$, which introduces stochasticity to the optimization. The algorithm for RED-DiffEq is shown in Algorithm 1.
Algorithm 1 RED-DiffEq for full waveform inversion.
1: Initialize velocity model: $\mathbf{x}_{0}$
2:for $k=0,\ldots,K-1$ do
3:  $t\sim\mathcal{U}\{1:T\}$ {Sample timestep uniformly (integer)}
9:end for
10:return $\mathbf{x}_{K}$ {Final velocity model}
The dominant cost of RED-DiffEq remains in the forward solver and differentiation process, as in conventional FWI. The diffusion-based regularization introduces a single U-Net inference per iteration, which is computationally negligible compared to wave-equation solves with a 0.8% runtime overhead (Sec. S4.5).